If person A takes 20 hours to complete the work and person B takes 30 hours to complete the same work, together in how much time the work will get done?
Answer by richwmiller(17219)

1/20+1/30=1/x
3/60+2/60=1/x
5/60=1/x
1/12=1/x
x=12